OpenStudy (anonymous):

-10(1/2x-1/5y)+30(1/6x+4/5y) how do I simplify this

OpenStudy (jdoe0001):

\(\bf -10\left(\cfrac{1}{2x}-\cfrac{1}{5y}\right)+30\left(\cfrac{1}{6x}-\cfrac{4}{5y}\right)\quad ?\)

OpenStudy (jdoe0001):

\(\bf -10\left(\cfrac{1}{2}x-\cfrac{1}{5}y\right)+30\left(\cfrac{1}{6}x-\cfrac{4}{5}y\right)\quad ?\)

OpenStudy (jdoe0001):

first off you'd need to distribute, of course

OpenStudy (anonymous):

Yes the 2nd one is correct

OpenStudy (jdoe0001):

\(\bf -10\left(\cfrac{1}{2}x-\cfrac{1}{5}y\right)+30\left(\cfrac{1}{6}x-\cfrac{4}{5}y\right)\implies {\color{red}{ -10}}\left(\cfrac{x}{2}-\cfrac{y}{5}\right)+{\color{red}{ 30}}\left(\cfrac{x}{6}-\cfrac{4y}{5}\right)\\ \quad \\ \left(\cfrac{{\color{red}{ -10}}x}{2}-\cfrac{{\color{red}{ -10}}y}{5}\right)+\left(\cfrac{{\color{red}{ 30}}x}{6}-\cfrac{{\color{red}{ 30}}\cdot 4y}{5}\right)\) see what you can cancel out from the denominator

OpenStudy (jdoe0001):

hmmm shoot ... got a bit truncated... ok so \(\bf -10\left(\cfrac{1}{2}x-\cfrac{1}{5}y\right)+30\left(\cfrac{1}{6}x-\cfrac{4}{5}y\right)\\ \quad \\\implies {\color{red}{ -10}}\left(\cfrac{x}{2}-\cfrac{y}{5}\right)+{\color{red}{ 30}}\left(\cfrac{x}{6}-\cfrac{4y}{5}\right)\\ \quad \\ \left(\cfrac{{\color{red}{ -10}}x}{2}-\cfrac{{\color{red}{ -10}}y}{5}\right)+\left(\cfrac{{\color{red}{ 30}}x}{6}-\cfrac{{\color{red}{ 30}}\cdot 4y}{5}\right)\)

OpenStudy (jdoe0001):

see any thing we can cancel out in the denominator?

OpenStudy (anonymous):

Ty for helping

OpenStudy (jdoe0001):

yw